OpenStudy (anonymous):

exact value of tan(2pi) so i turn this into a ratio sin(2pi)/cos(2pi) which goes to 0/1 = 0 . is this the right answer with the right reasons?

jimthompson5910 (jim_thompson5910):

yes, you are 100% correct

jimthompson5910 (jim_thompson5910):

you can also say that tan(2pi) = tan(0) tan(2pi)= sin(0)/cos(0) tan(2pi) = 0/1 tan(2pi) = 0 giving you the same thing